# Deep Learning for Economists

### v1

@@ Abstract
Deep learning provides powerful methods to impute structured information from large-scale, unstructured text and image datasets. For example, economists might wish to detect the presence of economic activity in satellite images, or to measure the topics or entities mentioned in social media, the congressional record, or firm filings. This review introduces deep neural networks, covering methods such as classifiers, regression models, generative AI, and embedding models. Applications include classification, document digitization, record linkage, and methods for data exploration in massive scale text and image corpora. When suitable methods are used, deep learning models can be cheap to tune and can scale affordably to problems involving millions or billions of data points.. The review is accompanied by a companion website, EconDL, with user-friendly demo notebooks, software resources, and a knowledge base that provides technical details and additional applications.
Deep neural networks have led to many recent scientific achievements – ranging from landing a rover on rugged Martian terrain to creating capable chatbots to transforming the diagnosis of disease. Deep neural networks typically map unstructured data - such as text, document image scans, satellite and other imagery, videos, and audio - to a continuous vector space. In the above examples, these vectors are used to compute the instructions to steer the spacecraft, autoregressively predict what word comes next given a prompt, or identify whether an image contains a tumor. Analogously, an economist might use neural networks to detect the presence of informal vendors in street view images, or to measure the topics or people mentioned in firm filings or government documents.
@@
Sequence Text Classify news 6.3
@@
Token Text Tag people, 6.4
@@
Paired text Text Text b 6.5
@@
Link Text, Link 7.2
@@
Link Text Link people 7.3
@@
Classification Text Track content 7.4
@@
Retrieval Images Optical character 7.5
@@
Object Images Detect document 8
@@
Table 1 summarizes this review’s applications. Most can be framed as classification problems, and the article also reviews regression, where a neural network is used to impute continuous values from text or images.
@@
The reader may be wondering how quickly this review will become outdated. It is helpful to consider the popular metaphor that neural networks are like Legos: different neural network components can be configured in various ways to achieve different ends, or to achieve a more state-of-the-art version of the same end. This review focuses on frameworks where it is straightforward to swap in new neural network components as the literature advances, e.g., replacing an older convolutional neural network with a vision transformer (Dosovitskiy et al., 2020), or updating a BERT language model backbone (Devlin et al., 2019) with the most recent language model. Technical and implementation details—those most likely to change as the literature advances—are provided on the accompanying EconDL website: https://econdl.github.io/. It provides a knowledge base organized into core topics, as well as links to open-source packages geared towards economists and pipelines that construct large-scale datasets with deep learning. Interested readers will find lecture notes and links to blog posts, textbook treatments, open courseware, and original papers. EconDL also links to demo notebooks for many of the applications in this review. The website will be updated on a continual basis, and some of the packages explicitly support swapping in new neural networks as the literature advances.
@@ 2.1 The Basics of Neural Networks
In a vanilla feedforward neural network, all neurons in one layer connect to all neurons in the next layer. Deep fully connected networks are rarely used in practice. Rather a few types of neural networks have dominated the deep learning literature. This review focuses on convolutional neural networks (CNNs; Section 2.2), recurrent neural networks (RNNs; Section 2.3), and the transformer (Sections 2.4, 2.5, and 2.6).
@@ 4 Bias and Uncertainty Quantification
There are many limitations to using deep learning to solve social or economic problems (see, for instance, papers from the ACM Conference on Fairness, Accountability, and Transparency https://facctconference.org/ and Cui and Athey (2022)). Here, our focus is much narrower: to impute or explore features of unstructured data that humans are likely to agree on, in contexts where the size of the raw dataset is orders of magnitude too large to extract features manually.
@@ 6.1 An Introduction to classifiers
Training a classifier is one of the most straightforward tasks in deep learning. The open-source package LinkTransformer can be used to train text sequence classifiers, with a demo available via EconDL. While the base transformer language model could be frozen when training a classifier, and various layers of the transformer could be used as inputs to the classifier, typically all parameters are allowed to update, with the classifier layer attached to the final layer of the transformer.
@@ 6.2 Generative AI for classification
This paper examines the performance of GPT-3.5 and GPT-4o on topic classification of historical newspaper articles as of June 2024. Over the past year, I have performed this exercise with older models, as well as GPT-4 and GPT-4 Turbo. GPT-4 and GPT-4o perform similarly, edging out Turbo. I haven’t seen any systematic improvements with new releases. Your mileage may vary, as there undoubtedly are tasks where the larger, newer (and hence more expensive) models will perform better.
I also examined two other leading AI models, from Anthropic: Claude Haiku and Claude Opus. These led to significantly worse performance (with F1 scores typically 10-40 points lower than GPT) and are not reported due to space constraints. The drivers of this lower performance are twofold. First, Claude refused to produce an output for texts that it assessed as harmful. A distinguishing feature of Claude is its ‘Constitutional AI’ framework, which sets out certain ethical principles (e.g., harmlessness necessitates that responses should be peaceful, ethical, and avoid content that might be considered offensive in non-Western cultures). Some articles on past conflicts (mostly objectively reporting on events) and a range of other topics (e.g., content about the introduction of contraception written in the 1960s) were considered harmful. Moreover, Claude didn’t always generate the desired Yes/No format, making it impossible to extract whether the article was on-topic. Neither of these behaviors arose with GPT. Perhaps this could be fixed with more prompt tuning, or it may change with a future update.
@@
The pros of generative AI for classification are: 1) startup costs are low, requiring minimal programming expertise or understanding of what is going on under the hood, and 2) it can be used zero-shot (without the user providing training data), whereas tuning a classifier requires training data. We will see that, if anything, custom-trained classifiers tend to have a performance edge on text classification, though the most straightforward tasks can be performed very well with generative AI. This is consistent with the broad consensus that human-level performance can generally be attained on supervised tasks given adequate, high-quality training data.
This brings us to potential disadvantages of generative AI. Using a large model behind an API does not provide the same fine-grained control as training a classifier. While models such as GPT allow users to expose the model to empirical examples, for the experiments in this paper, this did not lead to improvements in performance. It is not fully understood how demonstration through prompting conditions these models, whereas it is clear how providing training examples updates a classifier through gradient descent. Not all models learn equally when exposed to the same amount of training data (as demonstrated empirically in Section 9), and lightweight models like those used with customized classifiers tend to update very efficiently. A custom classifier also has an edge when it comes to interpretability and reproducibility. Results from a commercial API may no longer be reproducible if a model is deprecated, and as discussed above, commercial GenAI models can be more of a black box. These concerns could be mitigated by using an open-source foundation model, such as Meta AI’s Llama (Touvron et al., 2023). The startup costs and hardware requirements, however, negate the ease-of-use advantage. Finally, classifiers using a lightweight backbone such as RoBERTa (Liu et al., 2019) are cheap to deploy over a massive number of texts, whereas commercial models at present can be very expensive. This may change if competition increases and research on cheap deployment advances.
@@ 6.3 Sequence classification
Economists might wish to impute a variety of structured information at the level of a text: e.g., its topic, the type of content it contains, or its sentiment. To illustrate text sequence classification, this section trains 19 different binary topic classifiers, applied to massive-scale databases of historical news (Dell et al., 2023; Silcock et al., 2024), and compares them to generative AI. It would have been very difficult for annotators to keep 19 different topic definitions in mind to create multi-class labels; hence, binary classification is used. Binary classifiers cannot be combined into a multi-class classifier, as negatives for one topic may be positives for another. Keeping prompts simple for generative AI also suggests binary classes.
@@
A frequent question is how many labels are required. This will vary. Topics that are more diverse or require learning a more complex definition will require more labels. Topics that were seen many times in the pre-training of the language model may require fewer labels. Fortunately, training a classifier is compute-efficient. If, after the first round of training, results are unsatisfactory, it is straightforward to add more labels and retrain. An error analysis may give a sense of what types of texts require more training examples. Since labeling is costly, we recommend starting with fewer labels and adding more if needed.
Table 2 provides the split statistics for the various topic classification tasks examined in this section. [^2] The politics classifier is taken from a published paper (Dell et al., 2023) with different aims, and hence the split shares and overall number of annotations differ somewhat. Labeled data are randomly split into training, validation, and test data. Validation data are used to select hyperparameters, select the model checkpoint (when to stop training), and to tune prompts, whereas the test data were used only to compute Table 2. The prompts for Table 2 are listed in the supplemental materials.
@@
In most cases—across a diversity of topics—the tuned classifier tends to outperform or equal the performance of GPT, though for the more straightforward tasks, GPT’s performance can be very good, particularly in the case of GPT-4o (GPT-4 performed similarly). The training data used to produce these classifiers are high quality. With lower-quality labels, such as those created with online annotation platforms where quality is notoriously poor, a custom-trained classifier may well be consistently worse than GPT. These comparisons may also change in the future.
@@
At present, GPT is likely to be well beyond the budget of most social science researchers for large corpora, though I do not cite figures here as prices fluctuate and could change considerably depending on competition and technological advancement. In contrast, training a RoBERTa classifier on the number of labels shown here is very cheap (at the time of writing, it could be done within minutes on a \$9.99/month Google Colab plan or a mid-range Nvidia GPU card). I have also had students, with patience, train similar models on a laptop, though getting access to a decent GPU through cloud compute or dedicated hardware is preferable. Deploying the classifiers, even across millions of articles, is also cheap, and can be done either with cloud CPUs or on a mid-range GPU card in hours. One can circumvent the expense of generative AI by training a classifier on labels predicted by GPT. Table 2 shows this can work when GPT produces very high-quality labels. However, where GPT performs less well, training on noisy data can magnify errors.
@@
Figure 4: Share of newswire articles with a given binary topic tag.
@@ 6.4 Token classification
Figure 5: Shares of entity types in newswire articles.
@@ 7.3 Linking unstructured data
Entity mentions are disambiguated by embedding their contexts with the disambiguation model and retrieving their nearest Wikipedia neighbor in embedding space. If they are below a threshold cosine similarity to the nearest Wikipedia embedding, they are marked as not in the knowledge base. Figure 10 shows that entity mentions in 100 years of newswire coverage, spanning 1878-1977, correlate strongly with Wikipedia Qrank, which ranks Wikidata entities by aggregate page views in the Wikiverse Arora et al. (2024). Silcock et al. (2024) find that the most mentioned entity in these 100 years is Dwight Eisenhower—edging out Adolf Hitler, Richard Nixon, and Harry Truman—and only 4.7% of disambiguated entities in newswire articles are women.
Figure 10: Mentions against Wikipedia Qrank.
@@ 8 Regression
Regression is analogous to classification, except that a regression layer added to a neural network predicts a continuous number(s), rather than a set of class scores. For this reason, our treatment here is brief, focusing on a single application: object detection.

### v3

@@ Abstract
Deep learning provides powerful methods to impute structured information from large-scale, unstructured text and image datasets. For example, economists might wish to detect the presence of economic activity in satellite images, or to measure the topics or entities mentioned in social media, the congressional record, or firm filings. This review introduces deep neural networks, covering methods such as classifiers, regression models, generative AI, and embedding models. Applications include classification, document digitization, record linkage, and methods for data exploration in massive scale text and image corpora. When suitable methods are used, deep learning models can be cheap to tune and can scale affordably to problems involving millions or billions of data points.. The review is accompanied by a regularly updated companion website, EconDL, with user-friendly demo notebooks, software resources, and a knowledge base that provides technical details and additional applications.
1 Introduction
Deep neural networks have led to many recent scientific achievements – ranging from landing a rover on rugged Martian terrain to creating capable chatbots to transforming the diagnosis of disease. Deep neural networks typically map unstructured data - such as text, document image scans, satellite and other imagery, videos, and audio - to a continuous vector space. In other words, they map complex and diverse types of data into a format that is easier to process and understand. In the above examples, the resulting vectors are used to compute the instructions to steer the spacecraft, autoregressively predict what word comes next given a prompt, or identify whether an image contains a tumor. Analogously, an economist might use neural networks to detect the presence of informal vendors in street view images, or to measure the topics or people mentioned in firm filings or government documents.
@@ 1 Introduction
Sequence Text Classify news 7.3
@@
Token Text Tag people, 7.4
@@
Paired text Text Text b 7.5
@@
Link Text, Link 8.2
@@
Link Text Link people 8.3
@@
Classification Text Track content 8.4
@@
Retrieval Images Optical character 8.5
@@
Object Images Detect document 9
@@
Table 1 summarizes this review’s applications. Most can be framed as classification problems. The article also reviews regression, where a neural network is used to impute continuous values from text or images. It is important to note that the term “regression” is used somewhat differently in machine learning than it is in economics, which can generate confusion. In machine learning, “regression” refers to the prediction of a continuous number, whereas “classification” refers to the prediction of discrete outcomes. Regression encompasses any type of model used to predict continuous outcomes; in this review, we consider regression using deep neural networks.
@@
The reader may be wondering how quickly this review will become outdated. It is helpful to consider the popular metaphor that neural networks are like Legos: different neural network components can be configured in various ways to achieve different ends, or to achieve a more state-of-the-art version of the same end. This review focuses on frameworks where it is straightforward to swap in new neural network components as the literature advances, e.g., replacing an older convolutional neural network with a vision transformer (Dosovitskiy et al., 2020), or updating a BERT language model backbone (Devlin et al., 2019) with the most recent language model. Technical and implementation details—those most likely to change as the literature advances—are provided on the accompanying EconDL website: https://econdl.github.io/. It provides a knowledge base organized into core topics, as well as links to open-source packages geared towards economists and pipelines that construct large-scale datasets with deep learning. Interested readers will find lecture notes and links to blog posts, textbook treatments, open courseware, and original papers. EconDL also links to demo notebooks for many of the applications in this review. The website will be updated on a continual basis for as long as the transformer-based methods covered in this review remain state-of-the-art, and some of the packages explicitly support swapping in new neural networks as the literature advances.
@@ 3.1 The Basics of Neural Networks
In a vanilla feedforward neural network, all neurons in one layer connect to all neurons in the next layer. Deep fully connected networks are rarely used in practice. Rather a few types of neural networks have dominated the deep learning literature. This review focuses on convolutional neural networks (CNNs; Section 3.2), recurrent neural networks (RNNs; Section 3.3), and the transformer (Sections 3.4, 3.5, and 3.6).
@@ 5 Bias and Uncertainty Quantification
There are many limitations to using deep learning to solve social or economic problems (see, for instance, papers from the ACM Conference on Fairness, Accountability, and Transparency https://facctconference.org/ and Cui and Athey (2022)). Here, our focus is much narrower: to impute or explore low dimensional features of unstructured data that humans are likely to agree on, in contexts where the size of the raw dataset is orders of magnitude too large to extract features manually.
@@ 7.1 An Introduction to classifiers
Training a classifier is one of the most straightforward tasks in deep learning. The open-source package LinkTransformer can be used to train text sequence classifiers, with a demo available via EconDL. While the base transformer language model could be frozen when training a classifier, and various layers of the transformer could be used as inputs to the classifier, typically all parameters are allowed to update, with the classifier layer attached to the final layer of the transformer. Classifiers can be binary (two classes), multiclass (more than two classes, only one of which is positive), or multi-label (multiple classes can be positive).
@@
When using a classifier with more than two classes, the F1 score for each class can be calculated individually and then combined in different ways. One approach is Macro F1, where each class’s F1 score is averaged equally. Another method is Weighted F1, where each class’s F1 score is weighted by the number of true instances in that class. A third option is Micro F1, where global true positives, false positives, and false negatives are aggregated across all classes, and F1 is computed as in binary classification. Macro F1 is best suited when equal performance across classes is desired, while weighted or micro F1 are better choices when the dataset is imbalanced and performance on the more frequent classes is most important.
@@ 7.2 Generative AI for classification
This paper examines the performance of GPT-3.5 and GPT-4 on topic classification of historical newspaper articles. Over the past year, I have performed this exercise with older and newer models, as well as GPT-4o and GPT-4 Turbo. GPT-4 and GPT-4o perform similarly, edging out Turbo. GPT-4o mini performs worse. I haven’t seen any systematic improvements with new releases. Your mileage may vary, with the new releases that will no doubt be available by the time this article is published.
I also examined two other leading AI models, from Anthropic: Claude Haiku and Claude Opus. These led to significantly worse performance (with F1 scores typically 10-40 points lower than GPT) and are not reported due to space constraints. The drivers of this lower performance are twofold. First, Claude refused to produce an output for texts that it assessed as harmful. A distinguishing feature of Claude is its ‘Constitutional AI’ framework, which sets out certain ethical principles (e.g., harmlessness necessitates that responses should be peaceful, ethical, and avoid content that might be considered offensive in non-Western cultures). Some articles on past conflicts (mostly objectively reporting on events) and a range of other topics (e.g., content about the introduction of contraception written in the 1960s) were considered harmful. Moreover, Claude didn’t always generate the desired Yes/No format, making it impossible to extract whether the article was on-topic. Neither of these behaviors arose with GPT. Perhaps this could be fixed with more prompt tuning, or it could change with a new release.
@@
The pros of generative AI for classification are: 1) startup costs are low, requiring minimal programming expertise or understanding of what is going on under the hood, and 2) it can be used zero-shot (without the user providing training data), whereas tuning a classifier requires training data. We will see that, if anything, custom-trained classifiers tend to have a performance edge on text classification, but straightforward tasks can be performed well with generative AI. This is consistent with the broad consensus that human-level performance can generally be attained on supervised tasks given adequate, high-quality training data.
This brings us to potential disadvantages of generative AI. Using a large model behind an API does not provide the same fine-grained control as training a classifier. While models such as GPT allow users to expose the model to empirical examples, for the experiments in this paper, this did not lead to improvements in performance. It is not fully understood how demonstration through prompting conditions these models, whereas it is clear how providing training examples updates a classifier through gradient descent. Not all models learn equally when exposed to the same amount of training data (as demonstrated empirically in Section 10), and lightweight models like those used with customized classifiers tend to update very efficiently. A custom classifier also has an edge when it comes to interpretability and reproducibility. Results from a commercial API may no longer be reproducible if a model is deprecated, and as discussed above, commercial GenAI models can be more of a black box. These concerns could be mitigated by using an open-source foundation model, such as Meta AI’s Llama (Touvron et al., 2023). The startup costs and hardware requirements, however, negate the ease-of-use advantage. Finally, classifiers using a lightweight backbone such as RoBERTa (Liu et al., 2019) are very cheap to deploy over a massive number of texts, whereas commercial models at present can be expensive for large-scale problems. This may change if competition increases and research on cheap deployment advances.
@@ 7.3 Sequence classification
Economists might wish to impute a variety of structured information at the level of a text: e.g., its topic, the type of content it contains, or its sentiment. To illustrate text sequence classification, this section trains 19 different binary topic classifiers, applied to massive-scale databases of historical news (Dell et al., 2023; Silcock et al., 2024), and compares them to generative AI. It would have been very difficult for annotators to keep 19 different topic definitions in mind to create multi-class labels; hence, binary classification is used. Binary classifiers cannot be combined into a multi-label classifier, as negatives for one topic may be positives for another. Keeping prompts simple for generative AI also suggests binary classes.
@@
A frequent question is how many labels are required. This will vary. Topics that are more diverse or require learning a more complex definition will require more labels. Topics that were seen many times in the pre-training of the language model may require fewer labels. Fortunately, training a classifier is compute-efficient. If, after the first round of training, results are unsatisfactory, it is straightforward to add more labels and retrain. An error analysis may give a sense of what types of texts require more training examples. Since labeling is costly, I recommend starting with fewer labels and adding more if needed.
Table 2 provides the split statistics for the various topic classification tasks examined in this section. [^2] The politics classifier is taken from a published paper (Dell et al., 2023) with different aims, and hence the split shares and overall number of annotations differ somewhat. Labeled data are randomly split into training, validation, and test data. Validation data are used to select hyperparameters, select the model checkpoint (when to stop training), and to tune prompts, whereas the test data were used only to compute Table 2. The prompts for Table 2 are listed in the appendix (Table LABEL:tab:prompts-1.
@@
In most cases—across a diversity of topics—the tuned classifier tends to outperform or equal the performance of GPT, though for the more straightforward tasks, GPT’s performance can be very good, particularly in the case of GPT-4. The training data used to produce these classifiers are high quality. With lower-quality labels, such as those created with online annotation platforms where quality is notoriously poor, a custom-trained classifier may well be consistently worse than GPT. These comparisons may also change in the future with further innovation in open source and commercial models.
@@
Figure 4: Share of newswire articles with a given binary topic tag.
At present, GPT is likely to be beyond the budget of most social science researchers for massive corpora, though I do not cite figures here as prices fluctuate and could change considerably depending on competition and technological advancement. Training a RoBERTa classifier on the number of labels shown here is very cheap (at the time of writing, it could be done within minutes on a \$9.99/month Google Colab plan or a mid-range Nvidia GPU card). I have also had students, with patience, train similar models on a laptop, though getting access to a decent GPU through cloud compute or dedicated hardware is preferable. Deploying the classifiers, even across millions of articles, is also cheap, and can be done either with cloud CPUs or on a mid-range GPU card in hours. One can circumvent the expense of generative AI by training a classifier on labels predicted by GPT. Table 2 shows this can work when GPT produces very high-quality labels. However, where GPT performs less well, training on noisy data can magnify errors.
@@ 7.4 Token classification
Figure 5: Shares of entity types in newswire articles.
@@ 8.3 Linking unstructured data
Entity mentions are disambiguated by embedding their contexts with the disambiguation model and retrieving their nearest Wikipedia neighbor in embedding space. If they are below a threshold cosine similarity to the nearest Wikipedia embedding, they are marked as not in the knowledge base. Silcock et al. (2024) find that the most mentioned entity in these 100 years is Dwight Eisenhower—edging out Adolf Hitler, Richard Nixon, and Harry Truman—and only 4.7% of disambiguated entities in newswire articles are women.
@@ 9 Regression
In machine learning, the term “regression” refers to the prediction of continuous outcomes. Regression using deep neural networks is analogous to classification, except that a regression layer added to a neural network predicts a continuous number(s), rather than a set of class scores. For this reason, our treatment here is brief, focusing on a single application: object detection.